### v1

# Luth: Efficient French Specialization for Small Language Models and Cross-Lingual Transfer

Abstract
The landscape of Large Language Models (LLMs) remains predominantly English-centric, resulting in a significant performance gap for other major languages, such as French, especially in the context of Small Language Models (SLMs). Existing multilingual models demonstrate considerably lower performance in French compared to English, and research on efficient adaptation methods for French remains limited. To address this, we introduce Luth, a family of French-specialized SLMs: through targeted post-training on curated, high-quality French data, our models outperform all open-source counterparts of comparable size on multiple French benchmarks while retaining their original English capabilities. We further show that strategic model merging enhances performance in both languages, establishing Luth as a new state of the art for French SLMs and a robust baseline for future French-language research.
Luth: Efficient French Specialization for Small Language Models and Cross-Lingual Transfer
Maxence Lasbordes Université Paris-Dauphine Télécom SudParis maxence.lasbordes@dauphine.eu          Sinoué Gad École Polytechnique Télécom SudParis sinoue.gad@polytechnique.edu
1 Introduction
Large Language Models (LLMs) have shown great potential in complex multilingual tasks (Grattafiori et al., 2024; OpenAI, 2023; Yang et al., 2025), but performance is uneven across languages. Due to abundant English data, most research focuses on English, leaving other languages behind (Ruder et al., 2022; Li et al., 2024). French, spoken by over 280 million people, remains underrepresented in datasets and models, resulting in weaker performance within state-of-the-art multilingual systems.
In parallel, Small Language Models have emerged as a promising direction. Studies show that smaller models, when properly trained or adapted, can achieve competitive performance across diverse tasks (Lepagnol et al., 2024; Nguyen et al., 2024). Their compact size enables faster inference, lower computational overhead, and practical deployment, making them well-suited for real-world applications (Belcak et al., 2025). SLMs can also be efficiently specialized to specific languages or domains, offering a practical path to high-quality French language models without relying on large-scale resources.
2 Contributions
In this paper, we introduce Luth [^1] https://github.com/kurakurai/Luth, a family of compact French Small Language Models designed to address the English-centric bias through targeted adaptation. We demonstrate that using carefully curated post-training data, it is possible to significantly improve French capabilities, including general knowledge, instruction-following, and mathematical reasoning, without degrading original English performance, and even enhancing both languages through strategic model merging.
Specifically, our contributions are:
  1. 1.
    The Luth-SFT [^2] https://huggingface.co/datasets/kurakurai/luth-sft dataset, containing 570570k samples of French instruction-response pairs, which substantially improves model performance in general knowledge, instruction following, and mathematical reasoning.
  2. 2.
    The Luth [^3] https://huggingface.co/collections/kurakurai/luth-models-68d1645498905a2091887a71 family, comprising 5 models ranging from 350M to 1.7B parameters, achieving state-of-the-art performance in French within their size categories and delivering an absolute average improvement of up to $11.26\%$ across six French benchmarks.
  3. 3.
    An efficient and reproducible methodology for language-specific adaptation, easily extendable to other mid- and low-resource languages, while preserving performance in other languages.
Figure 1: Overview of the four main stages in constructing the Luth models, including their substeps, methods, and frameworks.
3 Related Work
The development of multilingual and language-specific models aims to mitigate the English-centric bias of current LLMs. Recent models such as BLOOM Le Scao et al. (2022), Llama Grattafiori et al. (2024) and AYA Üstün et al. (2024) cover dozens of under-represented languages, but they do not focus on language-specific optimization and frequently underperform on individual languages. Regional initiatives, for example EuroLLM [^4] EuroLLM and CroissantLLM fall within the category of Small Language Models, with parameter counts of 1.7B and 1.3B, respectively., target European languages (including French) to improve coverage and performance for that region Martins et al. (2024). Several efforts concentrate specifically on French. CroissantLLM Faysse et al. (2024) introduced a French–English bilingual model, Lucie Gouvert et al. (2025) has released and open-sourced a large portion of the resources needed for French LLM development, and Pensez Ha (2025) studied French model development with an emphasis on reasoning and argued for quality-over-quantity approaches to data curation. Despite these contributions, important gaps remain. Many works prioritize larger, resource-intensive models or report performance shortfalls relative to multilingual baselines of comparable size. They also provide few practical, low-cost recipes to substantially improve French-language capabilities, leaving room for work on compact, French-specialized models and efficient adaptation strategies suitable for resource-constrained deployments.
4 Luth-SFT Dataset
Figure 2: Overview of the Luth-SFT dataset construction pipeline, from data collection and translation to filtering and the Scholar subset creation.
To address the lack of high-quality open-source French post-training datasets, we introduce Luth-SFT, comprising 570570k samples ($338$ million tokens) of French instruction–response pairs (Figure 2).
Data Gathering  To construct this dataset, we first collected samples from existing multilingual datasets, including AYA Üstün et al. (2024), Smoltalk2, and CroissantLLM Faysse et al. (2024). These datasets contain texts in multiple languages; to efficiently isolate the French samples among hundreds of thousands of entries, we employed the langdetect library.
Data Translation  To further diversify and expand our French dataset, we selected two high-quality, openly available English instruction datasets, Tülu 3 Lambert et al. (2024) and OpenHermes. Our approach is twofold: (1) translate the English prompts into French (A.1) using strong multilingual models (GPT-4o and Qwen3 32B in non-reasoning mode), and (2) generate new French responses from scratch conditioned on the translated prompts, rather than directly translating the original answers. For Tülu 3, we focused exclusively on the math and instruction-following subsets, as these align with our objectives. The samples produced through this pipeline constitute the majority of our dataset. Notably, for OpenHermes, an existing French version generated with GPT-4o following this methodology was already available, substantially reducing the associated computational cost Alhajar (2025).
Filtering  We used a two-stage filtering pipeline to ensure both dataset quality and domain relevance. The first stage, linguistic validation, enforced strict French language criteria, including grammatical correctness, coherence, absence of code-switching or mixed-language content, and proper instructional formatting. The second stage, content filtering, systematically removed samples from three categories: programming-related content (e.g., code snippets, debugging queries, tool discussions), tool-calling content (e.g., API usage, command-line operations, system configuration), and samples containing logical inconsistencies or factual errors. This approach preserved instruction-following, scientific discourse, and general conversational samples while maintaining high linguistic and content standards. All system prompts used are listed in A.2.
Scholar  This subset was developed to address the scarcity of high-quality scientific resources in French. The dataset draws extensively from Baccalauréat and Classes Préparatoires aux Grandes Écoles (CPGE) examination materials, providing both questions and detailed solutions (see example snippet in A.3) across a broad range of subjects. A key objective was to build a resource that is non-synthetic and rooted in expert knowledge. Examination materials were particularly well-suited for this purpose, as they are typically accompanied by official solutions in PDF format, authored and validated by domain experts. In total, more than $\mathbf{14{,}000}\,\text{PDFs}$ were collected, covering examination sessions from 1980 to 2025555Mainly sourced from Sujet Bac and UPS Sujet.. These documents were processed through a multi-step pipeline (prompts listed in Appendix A):
  1. 1.
    Crawling and downloading the examination PDFs.
  2. 2.
    Filtering the documents (some PDFs contained scanned solutions and were therefore unusable).
  3. 3.
    Extracting questions and answers using a combination of regular expressions and LLM-assisted parsing with Gemini 2.5 Flash (Comanici et al., 2025).
  4. 4.
    Refining LaTeX formatting for equations and enriching the solutions with additional explanatory details (A.3) using Gemini 2.5 Pro (Comanici et al., 2025), as some official corrections were rather concise.
  5. 5.
    Performing a final filtering step to remove anomalous samples, including misaligned questions and answers, missing data, and formatting errors.
After processing, the dataset comprises $\mathbf{30{,}300~\text{samples}}$, distributed across several domains. The subject distribution is summarized Table 1. It should be noted that the proportions mainly reflect the availability of data for each subject, and do not represent a deliberate choice on our part.
Table 1: Distribution of scholars by subject.
Subject Percentage
Mathematics 67.23%
Physics–Chemistry 10.61%
Computer Science 9.08%
Engineering Science 6.04%
Biology 5.51%
Other (Economics, Accounting, Social Sciences) 1.52%
5 Luth Models
5.1 Model Training
As this work focuses on Small Language Models (SLMs) with fewer than 2B parameters, we conducted comprehensive evaluations of multilingual models in this size range to identify the best-performing model for French and to enhance its capabilities. We considered LFM2 (350M, 700M, and 1.2B) LiquidAI (2025) and Qwen3 (0.6B and 1.7B) Yang et al. (2025) for their strong French and English performance. While other SLMs, such as LLaMA 3.2 (1B) Grattafiori et al. (2024), SmolLM2 (360M and 1.7B) Allal et al. (2025), and Qwen2.5 (0.5B and 1.5B) Yang et al. (2024a), are also viable alternatives, our evaluations indicate that they underperform relative to more recent models on the tasks considered in this work. The models were selected based on their capabilities in Math, General Knowledge and Instruction Following in both French and English. Qwen3 and LFM2 variants then went through a full fine-tuning stage, instead of LoRA Hu et al. (2021) for better learning, on our Luth-SFT dataset, which infuses them with a richer understanding of French, specific vocabulary, domain-specific terminology, and improved their skills in the previously mentioned areas.
Full Fine-tuning  We fine-tuned the models on our curated Luth-SFT dataset using the Axolotl framework Axolotl maintainers and contributors (2023). The trainings were conducted on a single NVIDIA H100 GPU (80GB VRAM) for three epochs. We used various training hyperparameters for the models, which can be found in the Appendix B. For all models, we employed FlashAttention Dao et al. (2022) to reduce memory consumption and accelerate training through memory-efficient attention computation, and sequence packing to maximize GPU utilization by concatenating multiple shorter sequences into fixed-length batches, with a maximum sequence length of 16,38416{,}384. For instance, Luth-0.6B-Instruct was trained with widely used hyperparameters, including a learning rate of $2\times 10^{-5}$, an effective batch size of 24 (achieved via gradient accumulation), and a cosine learning rate scheduler with a $10\%$ warm-up period. Examples of training losses are shown in Figure 3. Due to computational limitations, we did not perform extensive hyperparameter sweeps for all models, and we leave this investigation to future work.
Figure 3: Loss per step during full fine-tuning on the Luth-SFT dataset over 3 epochs for Qwen3-0.6B (green) and Qwen3-1.7B (blue).
5.2 Model Merging
Model merging has recently gained attention as an effective technique for combining the parameters of multiple models, typically fine-tuned on different tasks or datasets, into a single system. This approach enables the merged model to inherit complementary strengths without additional retraining. Prior work has shown that merging can even outperform the individual components being merged (Yang et al., 2024b), a finding we confirm in our experiments (Figure 4).
Figure 4: Performance comparison of the Luth models in their base form (e.g., Qwen3-0.6B), after fine-tuning (e.g., Qwen3-0.6B fine-tuned), and after merging (e.g., Luth-0.6B-Instruct), averaged over four French/English benchmarks: IFEval, MMLU, GPQA-Diamond, and Math500. Left panel shows English performance, right panel shows French performance.
In our setting, this method is particularly relevant: since our dataset is exclusively French, fine-tuning strongly improves French capabilities but can slightly degrade performance in other languages, including English (Figure 4). Model merging offers a cost-effective solution to this problem, allowing us to preserve cross-lingual abilities while still gaining improvements in French. Indeed, we observe that merging not only recovers lost English performance but also improves overall results across both languages. Moreover, merging provides a natural way to mitigate catastrophic forgetting (Alexandrov et al., 2024).
Table 2: Overview of the Luth models and key merging details that produced the most stable performance across both French and English in our experiments. The coefficient (Coeff.) indicates the proportion of the fine-tuned model used in the merge with the base model (e.g., 0.7 corresponds to 70% of the fine-tuned model and 30% of the base model).
Model name Base model Merging method Coeff.
Luth-0.6B-Instruct Qwen3 SLERP 0.7
Luth-1.7B-Instruct Qwen3 SLERP 0.5
Luth-LFM2-350M LFM2 Linear 0.3
Luth-LFM2-700M LFM2 Linear 0.4
Luth-LFM2-1.2B LFM2 Linear 0.5
We used MergeKit, a framework that facilitates model fusion and provides a range of merging methods (Goddard et al., 2024). Since no single merging technique appears to be universally superior (Yang et al., 2024b), we experimented with various approaches. Surprisingly, the most stable results in our experiments were obtained with relatively simple methods, namely linear interpolation (LERP) and spherical linear interpolation (SLERP).
LERP combines two models in a straightforward linear fashion according to a coefficient $\alpha$:
$w=(1-\alpha)w_{0}+\alpha w_{1}$
SLERP, in contrast, interpolates along the arc of the unit sphere :
$w=\frac{\sin((1-\alpha)\theta)}{\sin(\theta)}w_{0}+\frac{\sin(\alpha\theta)}{\sin(\theta)}w_{1}$
with $\theta=\arccos(w_{0}\cdot w_{1})$, the angle between the two weights.
The main difference is that LERP follows a straight line in weight space, whereas SLERP follows a spherical arc, which can better preserve properties when the models are further apart.
We therefore empirically evaluated these methods and hyperparameters, and selected the ones that provided the best results, reported in Table 2.
6 Evaluation
Table 3: Results of Luth and other models on various French tasks. The scores are reported as percentages (Pass@1), averaged over three runs. The highest and second-best scores are shown in bold and underlined respectively for each model category.
Model
IFEval
French
GPQA-Diamond
French
MMLU
French
Math500
French
Arc-Challenge
French
Hellaswag
French
[…]
Qwen3-1.7B 54.71 31.98 28.49 60.40 33.28 24.86
SmolLM2-1.7B-Instruct 30.93 20.30 33.73 10.20 28.57 49.58
Qwen2.5-1.5B-Instruct 31.30 27.41 46.25 33.20 32.68 34.33
LFM2-1.2B 54.41 22.84 47.59 36.80 39.44 33.05
[…]
Llama-3.2-1B 27.79 25.38 25.49 15.80 29.34 25.09
LFM2-700M 41.96 20.81 43.70 32.40 36.27 41.51
Qwen3-0.6B 44.86 26.90 27.13 29.20 31.57 25.10
Qwen2.5-0.5B-Instruct 22.00 25.89 35.04 12.00 28.23 51.45
[…]
SmolLM2-360M-Instruct 21.50 28.43 26.14 3.20 26.60 32.94
LFM2-350M 31.55 28.93 38.63 18.00 33.36 39.13
To broaden our evaluation beyond French benchmarks, we also focused on English to assess our models’ capabilities. Our evaluation process is fully transparent, and all reported results are reproducible using open-source code [^6] https://github.com/kurakurai/Luth and publicly available data.
6.1 Benchmark Selection
As mentioned in the previous sections, we focused on specific capabilities in our training data, particularly instruction following, general knowledge, and mathematics. Among the dozens of English benchmarks available, we selected widely used ones that cover these specific capabilities. For French, we relied on benchmarks from multilingual efforts or on translated versions of their English counterparts, all openly available on Hugging Face. We used six benchmarks, each available in both French and English.
IFEval  IFEval Zhou et al. (2023) is a benchmark designed to evaluate instruction following and alignment abilities of language models, testing how well they adhere to and execute given instructions across diverse contexts.
Math500  Math Hendrycks et al. (2021b) is a mathematical reasoning dataset containing 500 problems ranging from arithmetic to higher-level mathematics, assessing models’ problem-solving and reasoning skills.
GPQA-Diamond  GPQA Rein et al. (2023) focuses on general knowledge question answering, providing challenging multiple-choice questions to test factual and commonsense reasoning.
MMLU  MMLU Hendrycks et al. (2021a) is a broad benchmark covering 57 subjects, including humanities and STEM, designed to evaluate general knowledge and multitask understanding.
Arc-Challenge  The AI2 reasoning challenge dataset Clark et al. (2018) consists of difficult multiple-choice science questions aimed at testing reasoning skills in grade-school science topics.
HellaSwag  HellaSwag Zellers et al. (2019) is a commonsense reasoning benchmark that requires models to select the most plausible continuation of a story or scenario, emphasizing context-dependent understanding.
6.2 Evaluation workflow
Most available evaluation frameworks provide limited support for French benchmarks, as they focus predominantly on English and offer minimal coverage of multilingual tasks. We chose to use LightEval Habib et al. (2024) due to its simplicity and its ability to easily add custom tasks. We added all the benchmarks mentioned above to our setup, along with their corresponding prompts and metrics in French.
The latest version of LightEval did not provide a mechanism to toggle reasoning mode for hybrid models. We modified it to add an enable_thinking option, allowing explicit control over the inclusion of reasoning traces enclosed in <think></think>. This extension was particularly important for Qwen3, which defaults to reasoning mode, as it enabled us to conduct all evaluations in non-reasoning mode.
We also extended LightEval to allow toggling enable_prefix_caching to false, since this feature is not supported by LFM2 models. Finally, we adapted the latest version of vLLM (0.10.2) to ensure compatibility with LightEval.
6.3 Results
We present the results of our five Luth models against several strong multilingual SLMs in Tables 3 and 4, for French and English respectively. Scores for each benchmark were computed as the average of three runs ($\text{temperature}=0$), using the same system prompts — "You are a helpful assistant." for English and "Vous êtes un assistant utile." for French.
Table 4: Results of Luth and other models on various English tasks. The scores are reported as percentages (Pass@1), averaged over three runs. The highest and second-best scores are shown in bold and underlined respectively for each model category.
Model
IFEval
English
GPQA-Diamond
English
MMLU
English
Math500
English
Arc-Challenge
English
Hellaswag
English
[…]
Qwen3-1.7B 68.88 31.82 52.82 71.20 36.18 46.98
SmolLM2-1.7B-Instruct 49.04 25.08 50.27 22.67 42.32 66.94
Qwen2.5-1.5B-Instruct 39.99 25.76 59.81 57.20 41.04 64.48
LFM2-1.2B 68.52 24.24 55.22 45.80 42.58 57.61
[…]
Llama-3.2-1B 44.05 25.25 31.02 26.40 34.30 55.84
LFM2-700M 65.06 30.81 50.65 32.00 38.65 52.54
Qwen3-0.6B 57.18 29.29 36.79 43.40 33.70 42.92
Qwen2.5-0.5B-Instruct 29.70 29.29 43.80 32.00 32.17 49.56
[…]
SmolLM2-360M-Instruct 33.95 20.71 26.18 3.00 35.41 52.17
LFM2-350M 56.81 27.27 44.79 20.87 34.27 45.07
Main insights  Luth models demonstrate that training on a high-quality, language-specific post-training dataset and leveraging model merging can lead to significant improvements in both French and English. Indeed, all Luth models substantially outperform their respective base models, as well as any model of comparable size, in French, while maintaining stable or even improved performance in English across widely used benchmarks. We attribute this phenomenon to cross-lingual transfer from French to English. Notably, Luth models exhibit average absolute score improvements in French ranging from $3.12\%$ to $11.26\%$ and in English from $0.76\%$ to $3.20\%$ across the six selected benchmarks. Furthermore, by fine-tuning the strongest SLMs available from two different families, we expect that our approach can substantially enhance the capabilities of any SLM under 2 billion parameters.
7 Conclusion
This paper introduces Luth, a family of state-of-the-art French Small Language Models (SLMs) that outperform all other models of comparable size on six French benchmarks covering general knowledge, instruction following, and mathematics. Although specialized in French, these models retain strong capabilities in other languages, particularly English, even showing improvements on various English benchmarks through cross-lingual transfer. These results stem from two key innovations: (1) the Luth-SFT, a French post-training dataset which drastically improves the model’s performance in French and (2) the use of model merging to retain multilingual skills while further improving each component’s specialized language capabilities. Moreover, we demonstrate that careful fine-tuning on a specific language alone can yield significant performance gains without resorting to costly methods like continual pretraining. We expect that similar improvements could extend to larger architectures and other languages; verifying this remains a direction for future work.
8 Limitations
While Luth models achieve state-of-the-art performance, several limitations remain. First, our evaluation covers only a limited set of benchmarks; while they provide strong signals, they do not fully capture the models’ capabilities.
Moreover, we assessed stability primarily in English, without thoroughly evaluating whether the models retain their abilities in other languages. Our experiments were also restricted to Small Language Models (under 2 billion parameters), which may limit the extent to which our approach unlocks potential gains at larger scales.
Finally, the Luth-SFT dataset does not cover key capabilities such as tool use or code generation, which are increasingly central to modern LLMs. While it significantly improves performance in French, it may not fully capture the language’s diverse structures and usage patterns.
Appendix A Luth-SFT System Prompts
A.1 Translation system prompt
System: You are a professional French translator. Translate English text into natural, accurate French. REQUIREMENTS: Preserve exact meaning, tone, and register of the original Use natural French syntax and idiomatic expressions Maintain all formatting (markdown, HTML, special characters, structure) Keep technical terms, code snippets, and proper nouns appropriately handled Ensure grammatical correctness and contemporary French usage OUTPUT: Only the translated French text in identical format to the input.
A.2 General filtering system prompts
System: You are a dataset quality assistant. Evaluate the question-answer pair below. Return False if any of the following apply: 1. Code-Related Programming questions or answers Code snippets or syntax Mentions of languages, libraries, tools Debugging or optimization 2. Tool Calling Content Describes or requests use of external tools, APIs, or systems Includes function calls, command-line usage, API requests, or tool invocation logic Involves configuring or troubleshooting external tools (e.g., databases, IDEs, browsers, CLIs) 3. Logical Errors Contradictions, invalid reasoning, factual errors 4. French Grammar Errors Wrong conjugations, tense, gender Bad structure, spelling, accents Respond only with: True or False
System: Validate French Q&A pairs. Return True only if BOTH the question and the answer meet all of the following criteria: Are written entirely in French. Are complete, grammatically correct, and coherent. Do not include any instruction to switch languages (e.g., “answer in English”, “répondez en anglais”, etc.). Do not contain mixed languages or foreign text (excluding proper nouns). The question must be an instruction or task prompt (e.g., “Traduis ce texte…”, “Explique…”). The question must not be a narrative, story, or purely informative content. Return False if any of these conditions are not met and respond with only: True or False.
A.3 Scholar
A.3.1 Extraction of Question/Answer pairs system prompt
System: The following is the full text of a French high school exam:  {subject}  For each question below, extract its introductory description from the subject (such as problem description or setup). Do not summarize or rewrite. Return a JSON list of dictionaries with keys: ‘‘question’’ and ‘‘context’’. Questions:
A.3.2 Refinement and enrichment system prompt
System: You will receive three inputs: a question, a context, and an answer. REQUIREMENTS: Correct any errors in spelling, grammar, LaTeX, and formatting in all three inputs. Carefully review the context and correct it if there are any issues. If the context is missing or empty but should be present based on the question and answer, generate a relevant and useful context. But do not provide the answer to the question or a hint. Rephrase the answer to add clarity by: Expanding on the reasoning, Breaking the answer down into logical steps or explanations, Justifying the conclusion. The final answer must remain logically and factually equivalent to the original. Do not change the overall intent of the question, context, or response unless necessary for correction. Provide your output strictly in JSON format with the following keys: ‘‘question’’, ‘‘context’’, and ‘‘reponse’’. The entire output must be in French. Do not add any hints in the context related to the question. ENTRIES: Question: {question} Context: {context} Response: {response}
A.3.3 Example snippet from the dataset
Role: User La suite $(u_{n})$ est définie pour tout entier naturel $n$ par $u_{n}=\int_{0}^{1}x^{n}\ln(1+x)\mathrm{d}x$.
Calculer $u_{0}$ à l’aide d’une intégration par parties.
Role: Assistant Pour calculer $u_{0}$, nous utilisons la définition de la suite $(u_{n})$ pour $n=0$. Étape 1 : Expression de $u_{0}$
En remplaçant $n$ par 0 dans l’expression générale, nous obtenons :
$u_{0}=\int_{0}^{1}x^{0}\ln(1+x)\mathrm{d}x=\int_{0}^{1}\ln(1+x)\mathrm{d}x$
Étape 2 : Application de l’intégration par parties
Nous calculons cette intégrale en utilisant la formule d’intégration par parties :
$\int_{a}^{b}f(x)g^{\prime}(x)\mathrm{d}x=[f(x)g(x)]_{a}^{b}-\int_{a}^{b}f^{\prime}(x)g(x)\mathrm{d}x$
Pour notre intégrale $\int_{0}^{1}\ln(1+x)\mathrm{d}x$, nous pouvons considérer que $\ln(1+x)$ est multiplié par 1. Nous posons alors : $f(x)=\ln(1+x)$, ce qui donne $f^{\prime}(x)=\frac{1}{1+x}$. $g^{\prime}(x)=1$. Pour la primitive $g(x)$, nous choisissons astucieusement $g(x)=x+1$. Étape 3 : Calcul de l’intégrale
En appliquant la formule avec nos fonctions, on obtient :
$u_{0}=[\ln(1+x)(x+1)]_{0}^{1}-\int_{0}^{1}\frac{1}{1+x}(x+1)\mathrm{d}x$
Simplifions et calculons chaque terme séparément : 1. Le terme entre crochets :
$(x+1)\ln(1+x)]_{0}^{1}=2\ln(2)-0=2\ln(2)$
2. L’intégrale restante :
$\int_{0}^{1}\frac{1}{1+x}(x+1)\mathrm{d}x=\int_{0}^{1}1\mathrm{d}x=1$
Étape 4 : Conclusion
$u_{0}=2\ln(2)-1$
Appendix B Training details
Table 5: Hyperparameters used to train Luth-0.6B-Instruct (Qwen3-0.6B) on a single Nvidia H100 80GB RAM.
Hyperparameter Value
Learning rate $2\times 10^{-5}$
Batch size (per device) 6
Gradient accumulation 4
Optimizer AdamW (8-bit)
Weight decay 0.01
Gradient clipping 0.1
Warmup steps 264
Scheduler Cosine
Max sequence length 16,384
Training epochs 3
Max training steps 2640
Precision bfloat16
Gradient checkpointing True
Flash Attention True
Packing True
Table 6: Hyperparameters used to train Luth-1.7B-Instruct (Qwen3-1.7B) on a single Nvidia H100 80GB RAM.
Hyperparameter Value
Learning rate $2\times 10^{-5}$
Batch size (per device) 3
Gradient accumulation 8
Optimizer AdamW (8-bit)
Weight decay 0.01
Gradient clipping 0.1
Warmup steps 264
Scheduler Cosine
Max sequence length 16,384
Training epochs 3
Max training steps 2640
Precision bfloat16
Gradient checkpointing True
Flash Attention True
Packing True
Table 7: Hyperparameters used to train Luth-LFM2-350M (LFM2-350M) on a single Nvidia H100 80GB RAM.
Hyperparameter Value
Learning rate $5\times 10^{-5}$
Batch size (per device) 8
Gradient accumulation 2
Optimizer AdamW (torch_fused)
Weight decay 0
Gradient clipping 0.1
Warmup steps 407
Scheduler Cosine
Max sequence length 16,384
Training epochs 3
Max training steps 4074
Precision bfloat16
Gradient checkpointing True
Flash Attention True
Packing True
Table 8: Hyperparameters used to train Luth-LFM2-700M (LFM2-700M) on a single Nvidia H100 80GB RAM.
Hyperparameter Value
Learning rate $5\times 10^{-5}$
Batch size (per device) 12
Gradient accumulation 3
Optimizer AdamW (torch_fused)
Weight decay 0.01
Gradient clipping 0.1
Warmup steps 270
Scheduler Cosine
Max sequence length 16,384
Training epochs 3
Max training steps 2709
Precision bfloat16
Gradient checkpointing True
Flash Attention True
Packing True
Table 9: Hyperparameters used to train Luth-LFM2-1.2B (LFM2-1.2B) on a single Nvidia H100 80GB RAM.
Hyperparameter Value
Learning rate $4\times 10^{-5}$
Batch size (per device) 8
Gradient accumulation 4
Optimizer AdamW (torch_fused)
Weight decay 0
Gradient clipping 0.1
Warmup steps 203
Scheduler Cosine
Max sequence length 16,384
Training epochs 3
Max training steps 2037
Precision bfloat16
Gradient checkpointing True
Flash Attention True
Packing True